### v2

# FinCon: A Synthesized LLM Multi-Agent System with Conceptual Verbal Reinforcement for Enhanced Financial Decision Making

Abstract
Large language models (LLMs) have demonstrated notable potential in conducting complex tasks and are increasingly utilized in various financial applications. However, high-quality sequential financial investment decision-making remains challenging. These tasks require multiple interactions with a volatile environment for every decision, demanding sufficient intelligence to maximize returns and manage risks. Although LLMs have been used to develop agent systems that surpass human teams and yield impressive investment returns, opportunities to enhance multi-sourced information synthesis and optimize decision-making outcomes through timely experience refinement remain unexplored. Here, we introduce the FinCon, an LLM-based multi-agent framework with conceptual verbal reinforcement tailored for diverse financial tasks. Inspired by effective real-world investment firm organizational structures, FinCon utilizes a manager-analyst communication hierarchy. This structure allows for synchronized cross-functional agent collaboration towards unified goals through natural language interactions and equips each agent with greater memory capacity than humans. Additionally, a risk-control component in FinCon enhances decision quality by episodically initiating a self-critiquing mechanism to update systematic investment beliefs. The conceptualized beliefs serve as verbal reinforcement for the future agent’s behavior and can be selectively propagated to the appropriate node that requires knowledge updates. This feature significantly improves performance while reducing unnecessary peer-to-peer communication costs. Moreover, FinCon demonstrates strong generalization capabilities in various financial tasks, including single stock trading and portfolio management.
1 Introduction
The intricacies and fluctuations of financial markets present substantial challenges for making high-quality, sequential investment decisions. In tasks such as single stock trading and portfolio management, each intelligent decision is informed by multiple market interactions and integrated information characterized by different levels of timeliness and modalities [1, 2]. These tasks aim to maximize profit while managing current market risks in an open-end environment.
In practice, trading firms often rely on synthesized human teamwork, with an organizational structure that typically involves hierarchical communication among various functional roles such as data analysts, risk analysts, portfolio managers, etc., and careful integration of various resources [3, 4]. However, the cognitive limitations of human team members can restrict their ability to quickly process market signals and achieve optimal investment results [5].
To improve investment returns and overcome human limitations, studies have employed methods such as deep reinforcement learning (DRL) to develop agent systems that simulate market environments and automate investment strategies [6, 7, 8]. Meanwhile, advancements in large language models (LLMs) demonstrate potential in complex tasks such as reasoning, tool-using, planning, and decision-making [9, 10], promising to surpass existing agent architectures. Language agents are known for human-like communication and adaptable prompt-based structures tailored for diverse decision-making settings [11, 12, 13, 14]. For optimal decision-making performance, two key elements need to be considered: 1) Organizing agents to promote effective teamwork and efficient communication, and 2) Enabling agents to continually learn and refine their actions. Studies indicate that imitating human organizational structures can effectively coordinate language agents for specific tasks [15, 16, 17]. Additionally, recent developments in textual gradient-based prompt optimization [18, 19] and verbal reinforcement [20, 21] have proven to be practical for enhancing the reasoning and decision-making abilities of language agents iteratively.
Language agent systems adapted for financial decision-making, such as FinGPT [22], FinMem [23] and FinAgent [24], have demonstrated impressive performance. However, these systems exhibit several limitations. Firstly, their reliance on the risk preferences of agents based on short-term market fluctuations fails to control overall risk exposures over long-term trading periods. This oversight can lead to potential losses due to the dismissal of fundamental factors driving investment profits. A more effective strategy is to quantify investment risks using measures of risk from quantitative finance [25, 26]. Secondly, most existing systems are designed for single-asset trading tasks and show limited adaptability to other financial applications involving multiple assets, such as portfolio management. Thirdly, these systems depend on the information understanding and extraction ability of a single agent, placing heavy pressure on the language model to comprehend and process information within a limited context window. This likely degrades decision quality. A multi-agent system that allocates tasks based on data source and modality could enhance performance. Even though approaches like StockAgent [27] utilize a multi-agent system for individual stock trading tasks. However, their decision-making processes heavily depend on extensive discussions among numerous LLM agents, leading to high communication costs and extended processing times. Besides, the lack of a clear optimization goal in such systems can compromise the effectiveness of the outcomes.
In response to these unaddressed issues, we propose FinCon, an LLM-based multi-agent framework for critical financial decision-making tasks such as single stock trading and portfolio management. Our main contributions are as follows: 1) Inspired by real-world investment roles, we designed an innovative Synthesized Manager-Analyst hierarchical communication structure and a risk control component. This communication structure distributes financial data from various sources and modalities to corresponding functional analyst agents, allowing them to focus on distilling essential investment insights and indicators from a single source of information. The manager then makes trading decisions by consolidating these insights and interacting with the risk-control component. The manager and analyst in a synchronous hierarchy to effectively enhance overall comprehension and reasoning capabilities. In this communication network, messages are conveyed only to the workers who need to be informed, eliminating redundant peer-to-peer communication and saving communication costs. 2) It is a generalized framework that can conduct not only stock trading but also portfolio management. The latter has not been claimed to be tackled by other financial language agent frameworks. 3) We innovatively crafted the agents’ persona using a dual-level risk control component, enabling comprehensive risk updates both within and across episodes. Within episodes, risk supervision is conducted using a quantile-based risk measure from quantitative finance, the Conditional Value at Risk (CVaR) [28]. For across-episode risk control, we designed a unique verbal reinforcement method to update the decision maker’s investment beliefs with conceptualized investment insights to guide future decisions. These investment beliefs are distilled from underlying reasoning trajectory-level profit and loss (PnL) changes and summarized as conceptual perspectives based on the functionality of all analyst agents. Investment beliefs can then be selectively back-propagated from the manager to the corresponding functional analyst agent. Ablation study results illustrate the significant effectiveness of our risk control design in managing market risk and improving trading performance.
2 Related Work
LLM Agents for Financial Decision Making. There are considerable efforts directed towards developing general-purpose LLM agent for sequential decision-making [29, 30], and such type of tasks often involve episodic interactions with environment and verbal reflections for action refinement, such as coding competition [31, 32], software development [33, 14], game-playing [34, 35]. Furthermore, researchers have started to exploit how LLM agents can perform better in harder decision-making tasks from finance [36, 37, 38], in which there are more volatile environments, leading to that the numerous unpredictable elements can obscure an agent’s ability to reflect accurately on the reasons for poor decision outcomes. FinMem [23] enhances single stock trading performance by embedding memory modules with LLM agent for reflection-refinement, and FinAgent [24] improved trading profits via using external quantitative tool to fight against volatile environment.
Multi-Agent System and Communication Structures. In traditional multi-agent systems [39, 40], the way for agents’ communication is pre-determined, like sharing data or state observations [41, 42, 43, 44, 45]. The emergence of large language model brings flexibility for human-understandable communications [46, 11, 14, 47], so some work tries to elevate decision-making ability of LLM-based multi-agent system by letting agents engage in discussions [48, 12] or debates [49, 50]. The similar peer-communication strategy was as well utilized by the multi-agent system for financial tasks [51, 52, 53]. However, such approach are not optimal for unified-goal financial tasks that prioritize profits [54], because they suffer from potentially ambiguous optimization objectives and are unable to control the unnecessary communication costs [55].
Prompt Optimization and Verbal Reinforcement. To enhance the reasoning or decision-making of LLM agents, many prompt optimization techniques have been proposed, like ReAct [56], Chain of Thought (CoT) [57], Tree of Thoughts (ToT) [58], ART [10], intended for that LLM agents can automatically generate intermediate reasoning steps as an iterative program. In addition, to make LLM agent make decisions like humans and generate more understandable reasoning texts, some researchers recommend incorporating cognitive structures [59, 60, 61, 62]. Inspired by these previous work and DRL algorithms [63], verbal reinforcement [20, 21, 64, 15] was developed for LLM agents such that they can update actions based on iterative self-reflection while integrating additional LLM as a prompt optimizer [18, 19].
3 Preliminaries
Here, we outline the mathematical notations for the two major financial decision-making tasks that will be explicitly discussed in our work. We also formally present the generalized modeling formulation using a Partially Observable Markov Decision Process (POMDP) [65] for financial decision-making tasks.
3.1 Financial Decision-making Tasks Formulation
Single Stock Trading Tasks. FinCon uses analyst agents group $\{M_{pr}^{i}\}_{i=1}^{I}$ to process multi-modal market information sources. The processed information is then used by a manager agent $M_{a}$ to make trading decisions (long, short, neutral), and to provide relevant reasoning texts. Note that the “short” signal means we make a “short-selling” decision, that is, a negative trading position is allowed. Additionally, FinCon evaluates the daily investment risk, followed by prompt-optimization for the manager agent from risk-control component $M_{r}$.
Portfolio Trading Tasks. In addition to processing multi-modal market information, the analyst agents also establish stocks’ pool for portfolio management, regarding statistical correlation between stocks’ returns. Then, the manager agent makes trading decisions on each stock in the pool. Lastly, the manager agent determines the portfolio weights for all stocks through a computational tool, which leverages the following mean-variance optimization [66]:
$\max_{\textbf{w}}\langle\textbf{w},\mu\rangle-\langle\textbf{w},\Sigma\textbf{w}\rangle\quad\text{s.t.}~{}w_{n}=\begin{cases}\in[0,1],&\text{``long''}\\ \in[-1,0],&\text{``short''}\\ =0,&\text{``neutral''}\end{cases},~{}~{}\forall n\in\{1,\cdots,N\}$
where $\textbf{w}=(w_{1}\cdots,w_{N})\in\mathbb{R}^{N}$ is portfolio weights vector, $\mu$ and $\Sigma$ are the shrinkage estimators of $N$-dimensional sample expected return and $N\times N$ sample covariance matrix of chosen stocks’ daily return sequences respectively [26]. We note that portfolio weights are rebalanced on daily basis. In our implementation, we first calculate the portfolio weights by solving the above optimization problem. Then, the target position is calculated by linearly scaling the portfolio weights from the prior step.
3.2 Modeling Quantitative Trading as POMDP
[…]
Then, our multi-agent system is supposed to learn the policies of all agents: the policies of analyst agents $\pi_{\theta^{i}}^{i}:\mathcal{X}\to\mathcal{A}^{i},i\in\{1,\cdots,I\}$ (the ways to process information, i.e. $\hat{O}_{t}^{i}\sim\pi_{\theta^{i}}^{i}(\cdot|O_{t}^{i})$), and the policy of manager agent $\pi_{\theta^{a}}:\mathcal{A}\times\mathcal{Y}\to\mathbb{A}$ (the ways to make trading decisions, i.e. $A_{t}\sim\pi_{\theta^{a}}(\cdot|\hat{O}_{t},B_{t})$) such that the system maximizes cumulative trading reward while controlling risk [70].
All policies $\Pi_{\bm{\theta}}=(\{\pi_{\theta^{i}}^{i}\}_{i=1}^{I},\pi_{\theta^{a}})$ are parameterized by prompts $\bm{\theta}=(\{\theta^{i}\}_{i=1}^{I},\theta^{a})$. By updating prompts via the risk-control component $M_{r}$, the whole system optimizes policies $\Pi_{\bm{\theta}}$ in a verbal reinforcement manner. By denoting daily profit & loss (PnL) by $R^{\Pi_{\bm{\theta}}}_{t}=\mathcal{R}(O_{t},B_{t},A_{t})$, the optimization objective for the whole system can be written as:
$\max_{\bm{\theta}}\mathbb{E}\Big{[}\sum\limits_{t\in\mathbb{T}}\alpha^{t}R^{\Pi^{\bm{\theta}}}_{t}\Big{]}$ (1)
which is a risk-sensitive optimization problem with respect to textual gradient descent, being essentially different from DRL algorithms for POMDP.
3.3 Textual Gradient-Descent
In an LLM-based prompt optimizer, a meta-prompt [18, 19] is used to refine the task prompt for better performance. For example, for a mathematical reasoning task, the task prompt might be "Let’s solve the problem," while the meta-prompt could be "Improve the prompt to help a model better perform mathematical reasoning."
Although prompt optimization lacks explicit gradients to control the update direction, we can simulate “textual gradient” by using LLMs’ reflection capabilities. By generating feedback from past successes and failures on trading decisions, LLMs can produce "semantic" gradient signals that guide the optimization process.
Adjusting the optimization process’s direction is crucial, similar to tuning the learning rate in traditional parameter optimization. An inappropriate learning rate can cause the process to oscillate or converge too slowly. Similarly, without proper control, the LLM-based optimizer might overshoot or oscillate during prompt optimization.
To mimic learning rate effects, we measure the overlapping percentage between trading decision sequences from consecutive iterations. We then directly edit the previous task prompt to enhance performance. The meta-prompt instructs the LLM to modify the current prompt based on feedback, ensuring a stable and incremental improvement process. This method allows for effective exploitation of existing prompts, leading to gradual performance enhancement.
Factor Gradient-based model optimizer LLM-based prompt optimizer Upgrade direction Model value gradient momentum Prompt reflection trajectory Update method Learning rate descent Overlapping percentage of trading decisions
Table 1: Analogy between glossaries in model optimizer and prompt optimizer.
4 General framework of FinCon
Figure 1: The general framework of FinCon.
In this section, we present the architecture of FinCon using a two-level hierarchy. First, we describe the hierarchical framework for coordinating the agents’ synchronous work and communication. Then, we elaborate on the functionalities of each module that constitutes each agent in FinCon.
4.1 Synthesized Multi-agent Hierarchical Structure Design
The agent system of FinCon consists of two main components: the Manager-Analyst Agent Group component and the Risk-Control component.
4.1.1 Manager-Analyst Agent Group
Analogous to human investment teamwork, FinCon establishes a unique hierarchy to organize the multi-agent system, synthesizing their efforts to achieve superior decision returns. The primary objective is to improve information presentation and comprehension while reducing unnecessary communication costs. The detailed working mechanism of each agent is explained and illustrated in Figure 2:
Figure 2: The detailed architecture of FinCon contains two key components: Manager-Analyst agent group and Risk Control. It also presents the between-component interaction of FinCon and decision-making flow.
Analyst Agents. In FinCon, analyst agents distill concise investment insights from large volumes of multi-source market information for one specific trading target. To ensure high reasoning quality by reducing task load and enhancing analysis focus, each agent processes a single information source in a uni-modal manner, delivering pre-specified outputs based on prompts. This setup mirrors an effective human team structure, where each analyst specializes in a specific function and synchronously filters market noise to extract essential insights. These agents support the manager agent in making informed decisions by consolidating denoised investment information from multiple perspectives. We design seven distinct types of analyst agents using LLMs, each producing diverse investment insights, as shown in the upper part of Figure 2. Differentiated by input modalities, three textual data processing agents distill insights and sentiments from daily news and financial reports. An audio agent extracts investment tendencies from earnings call recordings using the Whisper API. Additionally, a data analysis agent and a stock selection agent compute key financial metrics, such as momentum and CVaR, based on tabular time series data. The stock selection agent also manages portfolio selection following the classic risk diversification method in quantitative finance [1], as detailed in Appendix. 7.5.
Manager Agent. In the FinCon system, the manager agent is the sole decision maker, generating trading actions for sequential financial tasks. For portfolio management, it computes portfolio weights using convex optimization techniques constrained by the direction of actions (see Equation. (3.1)). Four mechanisms support each decision: 1) Consolidating distilled investment insights from multiple analyst agents. 2) Receiving timely risk alerts and conceptual investment updates from the risk control component. 3) Refining investment beliefs about the impact of various information sources on the manager’s decisions regarding a specific trading target. 4) Retrieving self-reflections based on the reasoning outcomes of previous trading actions.
4.1.2 Risk-Control Component
We innovatively design a dual-level risk-control mechanism comprising within-episode and over-episode risk management. The within-episode risk control mechanism detects market risk within a single training episode, enabling the manager agent to adjust trading actions promptly to prevent potential losses, which takes account of short-term trading performance and market fluctuations. Additionally, the within-episode risk control mechanism also operates during the testing stage. In contrast, the over-episode risk control mechanism only functions during the training stage, and it provides updating directions for prompt optimization by comparing the trading performance of the current episode with the previous one. This reflection helps the manager agent update its investment beliefs based on the differences in performance. Informed by prior observations of market risk and profitability patterns, these two mechanisms can potentially avoid repetitive investment mistakes, thereby enhancing gains for future investments.
Within-Episode Risk Control: . The within-episode risk alert is triggered by an instant drop in CVaR value. The rationale of this setting is that Conditional Value at Risk (CVaR) typically represents the average of the bottom 1% of past daily trading Profits and Losses (PnLs). When the CVaR value decreases on a given day, it often means that the PnLs from our trading decisions on the previous day (or days) are in the bottom 1%, indicating a potentially high-risk market condition. Once the within-episode risk is initiated, the manager agent is instructed to adopt a risk-averse attitude for the current day’s trading actions, regardless of the prior risk status.
Over-Episode Risk Control: The over-episode investment belief updates facilitate timely adjustments in the emphasis placed on the analyst’s information distillation and the manager’s action generation. Through the Actor-Critic mechanism, FinCon episodically optimizes its investment strategy for a given trading target by deeply reflecting on continuous winning and losing actions. The episodic reflection of investment beliefs is generated by a unique Conceptual Verbal Reinforcement (CVRF). CVRF evaluates the performance of every two adjacent training episodes based on the information perspectives provided by analysts and available in the manager’s trading reasoning. It then conceptualizes and attributes the evaluation outcomes to these aspects. By comparing the conceptualized aspects of investment insights from both more profitable and less profitable episodes, the system informs the manager and analyst agents about the necessary belief adjustments for each market information aspect to prioritize for greater profit, following the steps of Algorithm 1. CVRF utilizes text-based gradient descent to provide optimal conceptual investment guidance on a given target asset, substantially refining prompts with the latest investment beliefs for the manager agent. The obtained guidance is also meticulously organized from perspectives corresponding to insights from specific analyst agents, key financial indicators such as historical momentum, or other necessary perspectives.
These investment belief updates are initially received by the manager agent and selectively propagated to relevant agent nodes, thereby minimizing over-communication. Additionally, unlike the text-based gradient descent proposed by Tang et al.[19], which uses prompt editing distance as the learning rate, we obtain investment belief prompt updates by measuring the overlapping percentage of trading actions between two adjacent training trajectories achieved at each investment belief update. This approach has proven effective in enhancing the performance of a synthesized agent system, where each worker has a clearly defined and specialized responsibility.
Algorithm 1 Training Stage Algorithm of FinCon: Conceptual Verbal Reinforcement using Textual-based Gradient Descent
  Initialize manager-analysts component $\{M_{pr}^{i}\}^{I}_{i=1}\&M_{a}$, and risk-control component $M_{r}$.
  Initialize trading start date $s$, stock pool of portfolio and portfolio weights $w_{0}=\textbf{0}$.
  Initialize Prompts $\bm{\theta}$, policy $\Pi_{\bm{\theta}}$.
  while episode $k<Max$ do
     for $0\leq t\leq T$ do
        Run policy $\Pi_{\bm{\theta}}$ (collecting daily PnL $r_{t}$, portfolio weights $w_{t}$ and daily CVaR value $\rho_{t}$).
        if $\rho_{t}<\rho_{t-1}$ or $r_{t}<0$ then
           Trigger $M_{a}$ self-reflection and generate self-reflection text $B_{t}$.
        end if
        Get the investment trajectory $\mathcal{H}_{k}$ and calculate objective function value (1).
     end for
     Compare the objective function value 1 of episodes $k-1$ $k$, and decide which episode has higher performance;
     Pass sustained profitable and losing trades from two episodes $\mathcal{H}_{k-1}$ $\mathcal{H}_{k}$ into risk-control component $M_{r}$;
     Guide $M_{r}$ to summarize conceptualized investment insights $\{c_{k-1}^{1},\cdots,c_{k-1}^{n}\}$$\{c_{k}^{1},\cdots,c_{k}^{m}\}$;
     Compare two sets of conceptualized insights and give the reasoning for higher performance (providing textual optimization direction, i.e. $meta~{}prompt$);
     Calculate the overlapping percentage between trading decision sequences from two episodes (providing the learning rate $\tau$);
     Update the prompts by textual gradient-descent: $\bm{\theta}\longleftarrow M_{r}(\bm{\theta},\tau,meta~{}prompt).$
  end while
4.1.3 FinCon Test Stage Workflow
During the testing stage, FinCon will utilize the investment beliefs learned from the training stage, and the over-episode risk control mechanism will no longer operate. However, the within-episode risk control mechanism will still function, allowing the manager agent to adjust trading actions in real-time based on short-term trading performance and market fluctuations. This ensures that even during testing, FinCon can promptly respond to market risks and potentially prevent losses while leveraging the knowledge gained during training.
Algorithm 2 Testing Stage Algorithm of FinCon
  Initialize trading start date $s$, stock pool of portfolio and portfolio weights $w_{0}=\textbf{0}$.
  Inherit manager-analysts component $\{M_{pr}^{i}\}^{I}_{i=1}\&M_{a}$.
  Inherit the reflections $B$, learned prompts $\bm{\theta}$, the trained policy $\Pi_{\bm{\theta}}$.
  for $T+1\leq t\leq S$ do
     Run policy $\Pi_{\bm{\theta}}$ (collecting daily PnL $r_{t}$, portfolio weights $w_{t}$ and daily CVaR value $\rho_{t}$).
     if $\rho_{t}<\rho_{t-1}$ or $r_{t}<0$ then
        Trigger $M_{a}$ self-reflection.
     end if
     Get one investment trajectory $\mathcal{H}$.
  end for
  Output performance metrics calculation results based on $\mathcal{H}$.
4.2 Modular Design of FinCon Agents
Here, we explain the modular design of FinCon agents. Inspired by the recent works of Park et al. [61] and Sumers et al. [71] on developing the cognitive structure of language agents for human-like behavior, agents in FinCon integrate four modules to support their necessary functionalities, along with a shared general configuration, as detailed in Figure 3:
General Configuration and Profiling Module. The general configuration defines the task types (e.g., stock trading, portfolio management) and specifies the trading targets, including background information like trading sectors and performance overviews. The profiling module outlines the roles and responsibilities of each agent in text. The textual content from these two parts is concatenated and used to query investment-related events stored in the agents’ memory databases.
Perception Module. This module defines how an agent interacts with the market, specifying what they perceive, receive, and send. These aspects of information vary among different functional agents, as illustrated in Figure 3.
General Configuration 1. Investment task introduction    2. Trading target background 3. Trading sectors          4. Historical financial performance overview
Action Module Manager Agent:
1. Conduct: Trading actions.
2. Reflect: Trading reasons and analyst agents’ contribution assessment.
Profiling Module Manager Agent:
1. Role assignment: You are an experienced trading manager in the investment firm …
2. Role description: Your responsibilities are to consolidate investment insights from analysts and make trading actions on {asset symbols} … Analyst Agents:
1. Role assignment: You are the investment analysts for news/ market data/ Form 10-K (Q)/ ECC audio recording …
2. Role duty description: Your responsibilities are to distill investment insights and other indicators like financial sentiment for {asset symbols} …
Perception Module Manager Agent:
1. Perceive: Investment insights from analyst agents & Risk signal and trajectory-level investment belief updates from the risk control component.
2. Send: Feedback to analyst agents about their contribution to significant investment earnings & losses. Analyst Agents:
1. Perceive: Market information from certain information sources.
2. Receive: Feedback from the manager agent.
Memory Module Manager Agent:
1. Working:       - Consolidation          -Refinement
2. Procedural:    - Trading action records    - Reflection records 3. Episodic:       - Trajectory history Analyst Agents:
1. Working:        - Observation    - Retrieval    - Distillation
2. Procedural:     - Distilled Investment-related insights    - Financial sentiment - Investment report recommended actions
Figure 3: The detailed modular design of the manager and analyst agents. The general configuration and profiling modules generate text-based queries to retrieve investment-related information from the agents’ memory databases. The perceptual and memory modules interact with LLMs via prompts to extract key investment insights. The action module of the manager agent consolidates these insights to facilitate informed trading decisions.
Memory Module. The memory module contains three key components: working memory, procedural memory, and episodic memory. Similar to how humans process events in their working memory [72], FinCon agents utilize a working memory component to perform a series of operations, including observation, distillation, and refinement on available memory events, tailored to the agents’ specific roles. Procedural memory and episodic memory are essential for recording historical actions, outcomes, and reflections during sequential decision-making. Procedural memory is created after each decision step within an episode, storing data as memory events. For a trading inquiry, top events are retrieved from procedural memory, which are ranked based on recency, relevancy, and importance, which is the simplified version based on Yu et al. [23] and detailed in Appendix 7.6. Different functional analyst agents possesses distinct procedural memory decay rates, reflecting the timeliness of various financial data sources, which is crucial for aligning multi-type data impacting a specific time point and facilitating informed decision-making. The manager agent augments the procedural memory of analyst agents by providing feedback through access counter. Procedural memory is maintained by both analyst and manager agents, each keeping different records, as shown in Figure 3. The episodic memory, exclusive to the manager agent, retains actions, PnL series of the previous episodes and conceptualized updated investment beliefs from the risk control component.
5 Experiments
Reasearch Questions (RQs). Our experiment addresses three major research questions:
  • $\diamond$
    RQ1: Does FinCon demonstrate robustness across multiple financial decision-making tasks? We evaluate its performance on two challenging tasks: single-asset trading and portfolio management.
  • $\diamond$
    RQ2: Is the within-trajectory risk control mechanism in FinCon’s risk control component effective in maintaining superior decision-making performance?
  • $\diamond$
    RQ3: Is the over-trajectory risk control mechanism in FinCon’s risk control component effective for timely updating a manager agent’s beliefs and enhancing trading performance?
5.1 Experimental Setup
Multi-Modal Datasets. We collect a multi-modal dataset from real-world financial information to construct a representation of the market environment. It is composed of stock price information, daily financial news, company filing reports, and ECC audio recordings ranging from August 1, 2020, to August 14, 2023, as detailed in Appendix. 7.1. Each data source features distinct timeliness and is assigned to separate analyst agents to process. Typically, the time-sensitivity of annual filing reports (Form 10K’s) is more persistent, that of quarterly filing reports (Form 10Q’s) and ECC are secondary, and the daily financial news is the most instantaneous.
Evaluation Metrics. We evaluate the performance of FinCon and compare it with other state-of-the-art (SOTA) LLM-based and DRL-based agent systems using three standard financial performance metrics: Cumulative Return (CR%) [73], the Sharpe Ratio (SR) [74] and Max Drawdown (MDD%) [75]. Detailed definitions can be found in Appendix 7.3.
Comparative Methods. For the single stock trading task, we assess FinCon’s performance against seven algorithmic agents and the Buy-and-Hold (B & H) strategy, a widely accepted baseline. These include three DRL-based agents from the FinRL framework [76]: A2C, PPO, and DQN, and four SOTA LLM-based agents: Generative Agent by Park et al. [11], FinGPT [77], FinMem [23], and FinAgent [24]. For the portfolio management task, FinCon competes against the classical Markowitz MV portfolio selection strategy [1], the RL-based FinRL-A2C agent [76] and the B & H strategy that is to hold an equal-weighted position in all trading assets (so-called equal-weighted ETF). Our focus on classical portfolio strategy, RL methods and B & H methods is due to the limited availability of mature LLM agents for portfolio management. The detailed experiment parameter configuration is detailed articulated in Appendix. 7.7.
Implementation Details. In our experiments, all large language model (LLM)-based agent systems employ GPT-4-Turbo as their backbone algorithm, with the temperature parameter set at 0.3 to balance response consistency and reasoning creativity. FinCon is trained on financial data from January 3, 2022, to October 4, 2022, and is tested with data from October 5, 2022, to June 10, 2023. Given that deep reinforcement learning (DRL) algorithms require extensive training data for convergence, we extend the DRL agents’ training period to approximately five years, from January 1, 2018, to October 14, 2022, ensuring a fair comparison. The testing period remains consistent across all models. The performance metrics are reported for the test trajectory with the median CR and SR from five repeated epochs. (If the median CR and SR do not belong to the same epoch, the performance is based on the trajectory with the median CR.)
5.2 Main Results
In response to RQ1, we conduct analyses on two types of financial decision-making tasks: single-asset trading and portfolio management. The capability of FinCon to handle such sequentially complex decisions is thoroughly evaluated in the subsequent sections.
5.2.1 Single Asset Trading Task
In the first task, we evaluate the performance of FinCon against other leading algorithmic trading models by trading six different stocks. As shown in two tables above, FinCon significantly outperforms all the LLM-based and DRL-based approaches in terms of two primary metrics – cumulative returns and Sharpe Ratios. The FinCon’s Max Drawdown are also mostly among one of the lowest across all the trading assets, indicating a well-managed level of investment risk while still achieving the highest investment returns.
Overall, even with extended training periods, DRL-based models generally underperform. Specifically, the A2C algorithm significantly trails behind other algorithmic trading agents. Here, the training periods for Nio Inc (NIO) and Coinbase Global Inc (COIN) needs to be clarified. NIO completed its IPOs in September, 2018, so its training period is slightly shorter than other tickers. But the DRL algorithms for NIO still achieved convergence.The IPO for Coinbase Global, Inc (COIN) was completed in April 2021. With such limited trading data available, DRL algorithms struggle to converge, highlighting a significant limitation for DRL agents in automated trading of recent IPOs. Consequently, our analysis for COIN includes comparisons among FinCon, LLM-based trading agents, and the buy-and-hold strategy. In this analysis, FinCon demonstrates substantial performance advantages, achieving a cumulative return of over $57\%$ and a Sharpe ratio of $0.825$. We also note that LLM-based agents, capable of utilizing diverse data types and requiring minimal training, effectively address the challenges that DRL algorithms face.
Categories Models TSLA AMZN NIO
[…]
Market B&H 6.425 0.145 58.150 1.914 0.067 34.317 -77.210 -1.449 63.975
Our Model FinCon 82.871 1.972 29.727 24.964 0.906 25.889 17.461 0.335 40.647
LLM-based GA 16.535 0.391 54.131 -5.515 -0.195 37.213 -3.176 -1.574 3.155
FinGPT 1.549 0.044 42.400 -29.811 -1.805 -29.671 -4.959 -0.121 37.344
FinMem 34.624 1.552 15.674 -18.126 -0.776 36.825 -48.437 -1.180 64.144
FinAgent 11.960 0.271 55.734 -24.704 -1.496 33.151 0.933 0.051 19.181
DRL-based A2C -35.644 -0.805 61.502 -12.676 -0.447 37.179 -91.910 -1.728 68.911
PPO 1.409 0.032 49.740 3.863 0.137 28.085 -72.119 -1.352 62.093
DQN -1.296 -0.029 58.150 11.171 0.398 31.174 -35.419 -0.662 56.905
Categories Models APPL GOOG COIN
[…]
Market B&H 22.315 1.107 20.659 22.420 0.891 21.191 -21.756 -0.311 60.187
Our Model FinCon 27.352 1.597 15.266 25.077 1.0521 17.5299 57.045 0.825 42.679
LLM-based GA 5.694 0.372 14.161 -0.0151 -0.192 8.210 19.271 0.277 67.532
FinGPT 20.321 1.161 16.759 0.207 0.822 21.191 -99.553 -1.807 74.967
FinMem 12.396 0.994 11.268 0.311 0.018 21.503 0.811 0.017 50.390
FinAgent 20.757 1.041 19.896 -7.440 -1.024 10.360 -5.971 -0.106 56.882
DRL-based A2C 13.781 0.683 14.226 8.562 0.340 21.191 - - -
PPO 14.041 0.704 22.785 2.434 -0.097 25.202 - - -
DQN 21.125 1.048 16.131 20.690 0.822 21.191 - - -
Table 2: Comparison of key performance metrics during the testing period for the single-asset trading tasks involving six stocks, between FinCon and other algorithmic agents. Note that the highest and second highest CRs and SRs have been tested and found statistically significant using the Wilcoxon signed-rank test. The highest CRs and SRs are highlighted in red, while the second highest are marked in blue.
Moreover, in alignment with market trends, FinCon consistently exhibits superior sequential decision-making quality compared to other LLM-based financial agents, irrespective of whether the market conditions are bullish (e.g., GOOG), bearish (e.g., NIO), or mixed (e.g., TSLA). We attribute this performance to its high-quality distillation process through a synthesized multi-agent collaboration mechanism and its dual-functional risk control design, which collectively position FinCon as a leader in this space. Compared to FinCon, FinGPT primarily bases its trading decisions on the LLMs’ ability to interpret sentiment in financial information. This approach does not fully capitalize on the LLMs’ capacity to synthesize nuanced investment-related textual insights with numerical financial indicators. GA and FinMem, on the other hand, utilize single-agent frameworks that do not incorporate sophisticated information distillation processes or a diverse set of tools. This lack of complexity places considerable demand on the agent to process and reason through multi-sourced information, particularly when faced with large and varied data modalities. Additionally, they employ a static or minimal investment belief system. As a result, their ability to filter noise from mixed market data is relatively weak. As illustrated in Figure 7 (a) & (b), this limitation leads them to consistently maintain a lower position and exhibit hesitation regarding ‘buy’ or ‘sell’ decisions, ultimately resulting in suboptimal performance. FinCon overcomes these limitations through its innovative multi-agent synthesis, enabling it to deliver better outcomes. While FinAgent positions itself as a foundational framework capable of utilizing multi-modal data, its performance is robust with images and tabular data but does not exhibit a competitive edge when integrating audio data, such as ECC audio, which is crucial in real-world trading. Additionally, its decision-relevant memory retrieval process is based on similarity, leading the agent to make current decisions based on outdated memories, often resulting in errors. In contrast, the memory structure of FinCon is designed to account for the varying timeliness of multi-sourced financial data, ultimately delivering significantly superior decision quality.
Figure 4: CRs over time for single-asset trading tasks. FinCon outperformed other comparative strategies, achieving the highest CRs across all six stocks by the end of the testing period, regardless of market conditions.
5.2.2 Portfolio Management Task
To address the first research question, we compare FinCon’s performance with the Markowitz Mean-Variance (MV) portfolio [78] and FinRL [76] in managing a small portfolio of TSLA, MSFT, and PFE. Those assets are chosen by the stock selection agent from 42 stocks with sufficient amount of news data (over $800$ new messages in the training and test periods together), referring to Figure 9 in Appendix 7.2. The training & testing periods and backbone algorithm are consistent with those used in the single-asset trading task. For the Markowitz MV portfolio, we estimate the covariance matrix and expected returns using the same training data. For FinRL, we use 5-year training data before the test period. Our results, detailed in Table 3, show that FinCon outperforms both the Markowitz MV portfolio and FinRL, achieving significantly higher cumulative returns and Sharpe Ratio, and lower Maximum Drawdown (MDD).
However, there are stronger indications of hallucination in the multi-asset decision-making task compared to single-asset trading due to increased complexity and input length. Although FinCon reduces the workload of individual agents by distributing tasks and focusing on essential investment insights, it can still occasionally produce incorrect information, such as generating non-existent memory IDs and misclassifying action types (e.g., outputting ‘buy’ and ‘sell’ instead of ‘long’ and ‘short’). Handling multi-asset decision-making tasks requires intricate logic and substantial market information, posing a significant challenge for LLMs to comprehend extended contexts. This challenge has left portfolio management an untapped field for previous language agent studies. However, FinCon demonstrates the potential of constructing agent systems that can address complex financial problems through resource optimization, even when managing a relatively compact portfolio.
[…]
Table 3: Key performance metrics comparison among all portfolio management strategies. FinCon leads all performance metrics.
Figure 5: Portfolio value (CR for portfolio) changes over time for all the strategies.
5.3 Ablation Studies
In this section, we comprehensively evaluate our unique risk control component through two ablation studies. The first study assesses its ability to control risk within episodes using conditional value at risk (CVaR). The second study highlights the importance of the over-episode risk control mechanism in updating the trading manager agent’s beliefs for a holistic understanding of current trading circumstances. Consistency in training and testing periods is maintained with the main experiments for both single stock trading and portfolio management tasks.
5.3.1 The Effectiveness of Within-Episode Risk Control mechanism via CVaR
To answer the RQ2, we conduct the first ablation study. We assess the efficacy of FinCon’s within-episode risk control mechanisms by monitoring system risk changes through CVaR. To demonstrate the robustness of FinCon, we compare the performance of FinCon with versus without CVaR implementation across two task types: single-asset trading and portfolio management. Furthermore, in single-asset trading tasks, we consider assets in both general bullish and bearish market conditions in the testing phase for comprehensive consideration.
Our results demonstrate that implementing CVaR in FinCon is highly effective across all financial metrics for both task types, as shown in Table 4 and Fig 6. For single-asset trading tasks, FinCon without within-episode risk control yields negative CRs and significantly higher MDDs, underperforming compared to the Buy-and-Hold strategy (CR of GOOG: $22.42\%$, CR of NIO: $-51.057\%$), highlighting the severe consequences of ignoring environmental risks. In portfolio management, the CR increases dramatically from $16.994\%$ to $121.018\%$ with within-episode risk control, demonstrating its effectiveness in risk supervision even amid non-uniform market trends.
[…]
Table 4: Key metrics FinCon with vs. without implementing CVaR for within-episode risk control. The performance of FinCon with the implementation of CVaR won a leading performance in both single-asset trading and portfolio management tasks.
(a) [1] Single Stock: General Bullish
(b) [2] Single Stock: General Bearlish
(c) Multi-Assets
Figure 6: CRs of FinCon with vs. without implementing CVaR for within-episode risk control show that the CVaR mechanism significantly improves FinCon’s performance. This is evident from two metrics: (a) cumulative returns over time for single stocks in both bullish and bearish market conditions, and (b) portfolio value over time for a multi-asset portfolio. In both cases, FinCon with CVaR demonstrates substantially higher cumulative returns.
Specifically, the success of utilizing CVaR for within-episode risk control is evident in both bullish and bearish market environments, as shown in the single asset trading case. In bullish markets, CVaR sharply captures immediate market shocks and timely informs FinCon to exercise caution, even amidst general optimism. Conversely, in bearish markets, CVaR consistently alerts FinCon to significant price drops, ensuring awareness of market risks. Moreover, in portfolio trading with mixed price trends, our within-episode risk control mechanism performs robustly by monitoring the entire portfolio’s value fluctuations, enabling the trading manager agent to adjust potentially aggressive operations for each asset promptly.
5.3.2 The Efficacy of Over-Episode Belief Updates Using CVRF
In the second ablation study, to answer RQ3, we use the same assets to examine the effectiveness of FinCon’s over-episode risk control mechanisms. This is achieved by consistently improving FinCon’s beliefs about market conditions for the targeted assets. To ensure consistent belief output for each training episode, we set the temperature parameter to 0 specifically for belief generation.
We collect third-time belief updates over four training episodes using our innovative CVRF mechanism. The overlap of trading actions between the last two adjacent episodes increases to over $80\%$, and the updated investment beliefs are mostly aligned. To illustrate FinCon’s evolving investment beliefs through iterative training, we use the GOOG investment belief update as an example, as shown in Figure 8. Compared to the initial and final belief updates, each conceptual aspect, such as historical momentum and news insights, is enriched with executable information through our CRVF mechanism, leading to more profitable actions.
The results in Table 5 and Figure 7 indicate that the over-episode belief update mechanism is more critical than within-episode risk control in enhancing FinCon’s decision-making. Without this functionality, key metrics like CR, SR, and MDD are lower than without the within-episode risk control in single asset trading. Although the CR of $20.677\%$ slightly outperforms the Equal-Weighted ETF strategy’s $19.983\%$ for portfolio management, the SR and MDD are both lower. However, with the belief update feature, performance significantly improves, achieving a CR of $121.018\%$ and an SR of $3.435$. These results demonstrate that using CVRF to update investment beliefs over episodes efficiently steers the agent’s investment beliefs towards more profitable directions. FinCon achieves superior performance on multiple tasks, with learning gains evident after just four training episodes, requiring far fewer episodes than traditional RL algorithmic trading agents.
[…]
Table 5: Key metrics FinCon with vs. without implementing belief updates for over-trajectory risk control. The performance of FinCon with the implementation of CVaR won a leading performance in both single-asset trading and portfolio management tasks.
(a) [1] Single Stock: General Bullish
(b) [2] Single Stock: General Bearlish
(c) Multi-Assets
Figure 7: CRs of FinCon with vs. without belief updates for over-trajectory risk control. (a) The CRs over time for single stocks. The performance of FinCon with belief updates consistently leads in both bullish and bearish market conditions. (b) The CRs over time for multi-asset portfolio. FinCon’s performance with belief updates also won a substantially higher CR.
Updated Investment Belief Contents First Update: {’historical momentum’: ’Enhance the use of momentum indicators by establishing systematic rules for entry and exit based on momentum values to improve consistency and predictability in trading actions.’, ’news insights’: ’Integrate advanced real-time news sentiment analysis to better understand immediate market reactions and adjust trading strategies accordingly…’, ’Form 10-Q’: ’Incorporating annual report insights (10-K) could provide a comprehensive view of the company’s long-term performance and strategic direction, aiding in more informed decision-making.’…, ’other aspects’: [’sector trends’, ’technological advancements’], …} Last Update:
{’historical momentum’: ’The more profitable strategy effectively utilizes negative momentum values to make timely sell decisions, avoiding potential losses during downward trends, and effectively utilizes positive momentum values to make timely buy decisions, facilitating potential gains during upward trends. It is suggested to refine the use of momentum indicators, possibly by adjusting thresholds or incorporating additional short-term momentum metrics to enhance the timing and accuracy of buy/sell decisions.’,
’news insights’: ’It is recommended to further develop a systematic approach to quantify the impact of news, especially the sentiment scores and regulatory changes, to refine trading decisions.’, ’Form 10-Q’: ’It is suggested that even if there is a strong financial performance present 10-Q reports, it is better to prioritize immediate market signals. For future strategies, it is suggested to balance these aspects more evenly, especially in stable or bullish market conditions, to avoid premature exits from potentially profitable positions.’,…, ’other aspects’: [’sector trends’, ’technological advancements’, ’macroeconomic factors’, ’regulatory risks’], …}
Tradng Action Overlapping Percentages 1. Second vs. First Training Episode:   $46.939\%$ 2. Third vs. Second Training Episode: $71.429\%$ 3. Fourth vs. Third Training Episode:  $81.633\%$
Figure 8: The first time and last time LLM generated investment belief updates by CVRF for GOOG.
6 Conclusion
In this paper, we present FinCon, a novel LLM-based multi-agent framework for financial decision-making tasks, including single stock trading and portfolio management. Central to FinCon is the Synthesized Manager-Analyst hierarchical communication structure and a dual-level risk control component. This communication method channels financial data from multiple sources to specialized analyst agents, who distill it into key investment insights. The manager agent then synthesizes these insights for decision-making. Our experimental evaluations demonstrate the efficacy of our risk control mechanism in mitigating investment risks and enhancing trading performance. Additionally, the streamlined communication structure reduces overhead. The dual-layered risk control component introduces a novel approach to defining agent personas, enabling dynamic updates of systematic risk and market beliefs within agent communication. A valuable future research direction would be to scale FinCon’s framework to manage large-sized portfolios comprising tens of assets, while maintaining the impressive decision-making quality demonstrated with smaller portfolios. Given the LLM’s input length constraint, a critical challenge lies in striking an optimal balance between information conciseness through agent distillation and potential performance deterioration when extending the current context window. Addressing this will be essential for ensuring quality-assured outcomes.